# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ 2.4 Phase-Wise Modeling of the Energy Equation
Note that instead of pursuing a one-fluid formulation for the energy equation which includes additional source terms [11, 12, 13, 35] like it is done for the momentum equation, Eq. (16) is solved separately in a phase-wise manner in each phase with the interface acting as a moving boundary [8, 38, 16, 17]. The thermodynamic equilibrium at the interface defines an interface temperature (e.g., saturation temperature), which is embedded in the discretization of convective and diffusive terms. Generally, the interface temperature can vary depending on the local equilibrium conditions. In this work, the analysis is limited to saturated conditions with a uniform interface temperature. The interface state is then implicitly accounted for; thus, no additional source terms are required. The phase-wise approach allows for a sharp treatment of the energy jump condition and a reduction of the smearing of the temperature field compared to the one-fluid formulation. Albeit out of scope, a similar strategy can be used to solve the species transport equations in multi-component multiphase flows.
@@ 3 Flow Solver Algorithm
    Calculate new $\dot{m}^{\prime}$ based on $T^{n+1}$ and $C^{n+1}$
@@
    Calculate $\mathbf{f}_{\sigma}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ based on new $\dot{m}^{\prime}$ and $C^{n+1}$
@@
    Calculate $\mathbf{f}_{NC}$ based on new $\dot{m}^{\prime}$, $C^{n+1}$ and $\mathbf{u}^{*}$
@@ 3.2 Calculation of Interfacial Mass Flux
Different approaches have been implemented in the literature to obtain $\dot{m}^{\prime}$ from Eq. (8), focusing on the calculation of the temperature gradients in each phase, e.g., [8, 38, 10, 42, 13, 1, 17, 35]. Here, a normal stencil or probe is extended from the interface plane centroid into each phase along $\mathbf{n}_{\Gamma}$ [43, 44, 16]. Then, scalar quantities from the grid nodes, i.e., temperature, are linearly interpolated on two probing points on each side. To avoid using values from the opposite phase during interpolation, the first point has to be placed sufficiently far away from the interface. One can determine that the minimum distance to avoid this conflict is the cell diagonal (i.e., $\sqrt{2}\Delta x$ in two dimensions and $\sqrt{3}\Delta x$ in three dimensions). Consequently, the first point is placed at a distance of $\Delta x^{(1)}=1.75\Delta x$ from the interface. Additionally, a second point is placed at $\Delta x^{(2)}=\Delta x$ from the first point (i.e., $\Delta x^{(1)}+\Delta x^{(2)}=2.75\Delta x$). A two-dimensional sketch is shown in Figure 2.
One-sided second-order finite differences with non-uniform spacing can be used to calculate
@@
where, oftentimes, $T_{\Gamma}$ equals the saturation temperature at the given system pressure $T_{sat}$. An inherent issue of the calculation of $\nabla T_{G}\cdot\mathbf{n}_{\Gamma}$ and $\nabla T_{L}\cdot\mathbf{n}_{\Gamma}$ is its effect on the smoothness of the $\dot{m}^{\prime}$ distribution. As such, some works perform weighted averages of the temperature gradients around the interface cell to obtain a more uniform distribution [10, 13, 17]. Despite a smoother $\dot{m}^{\prime}$ distribution is observed with the proposed gradient calculation compared to other works for similar grid resolutions, variations in $\dot{m}^{\prime}$ still affect the long-term evolution of canonical configurations such as static evaporating droplets or growing bubbles. Therefore, once Eq. (8) is evaluated, the value of $\dot{m}^{\prime}$ is recalculated at each interface cell by averaging $\dot{m}^{\prime}$ from interface cells contained in a stencil of size $3\times 3\times 3$. As shown in Section 5.1, even though the first probing point extends almost two grid sizes into each phase, the second-order one-sided scheme results in $\dot{m}^{\prime}$ errors and convergence rates similar to other works.
@@ 3.3 Solution of the Navier-Stokes Equations
The Navier-Stokes equations are solved using a modification of the predictor-projection method [45] together with direct solvers based on Fast Fourier Transforms (FFT) for Poisson-type equations [46, 47], which have been integrated into the solver with the vfftpack library [^2] A vectorized package of Fortran subprograms for the fast Fourier transform of multiple real sequences, R.A. Sweet, L.L. Lindgren and R.F. Boisvert, https://www.netlib.org/vfftpack/.
@@
Then, Eq. (21) is a Poisson equation for $\psi$ obtained by taking the divergence of Eq. (20), which is solved using the FFT-based or direct solver with the same boundary conditions as the pressure field. As shown later in Section 3.5, a pressure jump would be introduced if this first predictor-projection step were not included (i.e., $\mathbf{u}^{*}=\mathbf{u}^{n}$), resulting from the different Stefan flow from $t^{n}$ to $t^{n+1}$ and directly linked to the first term on the LHS of Eq. (12).
$\nabla^{2}\psi=\nabla\cdot\mathbf{u}^{n}-\nabla\cdot\mathbf{u}^{*}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n}-\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n+1}$ (21)
@@
with, e.g., $\nabla\psi_{i+\frac{1}{2}}\approx(\psi_{i+1}-\psi_{i})/\Delta x$ at the cell faces. Note some indices have been dropped for convenience. Here, $\delta_{\Gamma}=||\nabla C||$ is evaluated from staggered volume fractions to minimize the smearing of the velocity jump when compared to using central differences. For example, Figure 1 shows how $C_{i+\frac{1}{2}}$ is obtained from adding the enclosed volumes on the right half of $C_{i}$ and on the left half of $C_{i+1}$. This evaluation of $\delta_{\Gamma}$ produces a smoother velocity distribution around curved interfaces compared to obtaining $\delta_{\Gamma}$ from the surface area density, e.g., [10, 16].
Next, the second predictor step calculates a velocity $\mathbf{u}^{**}$ from Eq. (15) without the pressure gradient, $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$ using the AB scheme
@@
As done for the integration of the energy equation, AB is used everywhere outside $\Omega_{\Gamma}$, while a classic first-order scheme is used in $\Omega_{\Gamma}$. The viscous term is discretized with second-order central differences, and the convective term is discretized with the third-order WENO scheme in cells not belonging to $\Omega_{\Gamma}$ and second-order central differences in $\Omega_{\Gamma}$. These modifications around the interface are necessary to aim for the cancellation of the numerical jump caused by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ with $\mathbf{f}_{NC}$. Note the choice of WENO in the momentum equation is to maintain consistency with the discretization of the energy equation. For practical purposes, no major differences have been observed between using WENO or central differences in the problems considered in this work.
The variable-coefficient pressure Poisson equation (PPE) resulting from the projection step is converted into a constant-coefficient PPE via the approximate substitution given by the split pressure gradient method [48, 9]
@@
which separates $\tfrac{1}{\rho^{n+1}}\nabla p^{n+1}$ into a constant-coefficient implicit term and a variable-coefficient explicit term. For accuracy and stability purposes, $\rho_{0}=\text{min}(\rho_{g},\rho_{l})$ and $\hat{p}=[1+(\Delta t^{n+1}/\Delta t^{n})]p^{n}-(\Delta t^{n+1}/\Delta t^{n})p^{n-1}$ is a linear extrapolation in time of the pressure field [48, 9]. Then, a direct pressure solver can be used, resulting in a computationally more efficient solver than iterative methods [9, 49, 11, 12, 16].
However, pressure-splitting introduces considerable time-stepping restrictions for stability purposes when the pressure jump across the interface is large (e.g., large surface tension) or when the density ratio is large (i.e., $\nabla\hat{p}$ has a greater contribution in the PPE) [9]. In other words, smoothness in $\hat{p}$ may be lost and substantial pressure oscillations occur as the interface moves across grid cells [50, 51]. Such limitation could defy the implementation of the gradient splitting technique and reduce the benefits of a direct solver. This problem is mitigated by explicitly accounting for the pressure jump in the PPE [50], similar to a GFM. Thus, $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$ are removed from Eq. (15) during the predictor step in Eqs. (23) and (24) and, instead, are added in the projection step. Within the context of the gradient splitting technique, this is equivalent to the temporal extrapolation of the pressure jump instead of the pressure itself, i.e., see Eq. (26).
@@
$\mathbf{u}^{n+1}=\mathbf{u}^{**}-\Delta t^{n+1}\Bigg{[}\frac{1}{\rho_{0}}\bigg{(}\nabla p^{n+1}-\mathbf{f}_{\sigma}^{n+1}-\mathbf{f}_{\dot{m}^{\prime}}^{n+1}\bigg{)}+\bigg{(}\frac{1}{\rho^{n+1}}-\frac{1}{\rho_{0}}\bigg{)}\bigg{(}\nabla\hat{p}-\hat{\mathbf{f}}_{\sigma}-\hat{\mathbf{f}}_{\dot{m}^{\prime}}\bigg{)}-\frac{1}{\rho^{n+1}}\mathbf{f}_{NC}^{n+1}\Bigg{]}$ (26)
@@
$\nabla^{2}p^{n+1}=\nabla\cdot\bigg{[}\bigg{(}1-\frac{\rho_{0}}{\rho^{n+1}}\bigg{)}\bigg{(}\nabla\hat{p}-\hat{\mathbf{f}}_{\sigma}-\hat{\mathbf{f}}_{\dot{m}^{\prime}}\bigg{)}+\mathbf{f}_{\sigma}^{n+1}+\mathbf{f}_{\dot{m}^{\prime}}^{n+1}+\frac{\rho_{0}}{\rho^{n+1}}\mathbf{f}_{NC}^{n+1}\bigg{]}+\frac{\rho_{0}}{\Delta t^{n+1}}\big{[}\nabla\cdot\mathbf{u}^{**}-\nabla\cdot\mathbf{u}^{n+1}\big{]}$ (27)
@@
where $\nabla\cdot\mathbf{u}^{n+1}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n+1}$ ensures that the correct velocity jump given by Eq. (10) is imposed. It is important to highlight that the splitting operator is not applied to $\mathbf{f}_{NC}^{n+1}$ since the variable density $\rho^{n+1}$ effectively cancels with the density term inside $\mathbf{f}_{NC}^{n+1}$. Then, Eq. (27) is solved using the FFT-based or direct solver. Another advantage of using the direct solver relates to stiffness problems that might arise as $\nabla\cdot\mathbf{u}^{n+1}$ increases, which could affect the convergence rate of iterative solvers [8, 11].
@@
The forces $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$ must be calculated at cell faces to correspond with the evaluation of $\nabla p$. In $\mathbf{f}_{\sigma}$ and $\mathbf{f}_{\dot{m}^{\prime}}$, $\kappa$ and $\dot{m}^{\prime}$ are averaged at cell faces [9] and $\mathbf{n}_{\Gamma}\delta_{\Gamma}=-\nabla C$ following the CSF approach. $\nabla C$ is evaluated from cell values, e.g., $\partial C/\partial x|_{i+1/2}\approx(C_{i+1}-C_{i})/\Delta x$. In contrast, $\mathbf{f}_{NC}$ is discretized differently for smoothness purposes in order to minimize geometrical errors that may arise from an ill-defined discrete force at the interface given the sensitivity of the term $\mathbf{u}^{*}\cdot\mathbf{n}_{\Gamma}$. Note that the corrective nature of $\mathbf{f}_{NC}$ to counteract the error introduced by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ in $\Omega_{\Gamma}$ implies that $\mathbf{u}^{*}$ substitutes $\mathbf{u}$ in Eq. (13). In this work, $\mathbf{u}^{*}$ is linearly averaged at cell faces while $\mathbf{n}_{\Gamma}$ is obtained at staggered interface cells (e.g., $10<C_{i+1/2}<$ 1) using the MYC method with staggered volume fractions. Then, a narrow band of staggered cells around the interface is populated with a weighted average of the normal unit vectors obtained in the previous step with the method described in A. Lastly, discretizing $\mathbf{n}_{\Gamma}\delta_{\Gamma}$ with $\delta_{\Gamma}$ obtained in staggered locations as, e.g., $\delta_{\Gamma,i+1/2}=\tfrac{1}{2}(\delta_{\Gamma,i}+\delta_{\Gamma,i+1})$ improves the smoothness of $\mathbf{f}_{NC}$ compared to the substitution $\mathbf{n}_{\Gamma}\delta_{\Gamma}=-\nabla C$. As shown in Section 3.5, both $\delta_{\Gamma,i+1/2}$ or $\nabla C$ evaluated with staggered central differences, e.g., $\partial C/\partial x|_{i+1/2}\approx(C_{i+3/2}-C_{i-1/2})/(2\Delta x)$, result in the correction of the error introduced by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ at the discrete level for a well-posed one-dimensional case.

### v2

@@ 2.4 Phase-Wise Modeling of the Energy Equation
Instead of pursuing a one-fluid formulation for the energy equation which includes additional source terms [11, 12, 13, 14, 35], Eq. (16) is solved separately in a phase-wise manner in each phase with the interface acting as a moving boundary [8, 38, 16, 17]. Both options are commonly used in the literature, but in the latter the thermodynamic equilibrium at the interface defines an interface temperature (e.g., saturation temperature), which is embedded in the discretization of convective and diffusive terms. Generally, the interface temperature can vary depending on the local equilibrium conditions. In this work, the analysis is limited to saturated conditions with a uniform interface temperature. The interface state is then implicitly accounted for; thus, no additional source terms are required. The phase-wise approach allows for a sharp treatment of the energy jump condition and a reduction of the smearing of the temperature field compared to the one-fluid formulation. Albeit out of scope, a similar strategy can be used to solve the species transport equations in multi-component multiphase flows. Note that a one-fluid formulation of the energy equation could be developed similar to the proposed momentum equation to ensure that the energy balance is satisfied [14]. However, the sharpness of the phase-wise approach has been favored in this work.
@@ 3 Flow Solver Algorithm
    Calculate new $\dot{m}^{\prime\prime}$ based on $T^{n+1}$ and $C^{n+1}$
@@
    Calculate $\mathbf{f}_{\sigma}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ based on new $\dot{m}^{\prime\prime}$ and $C^{n+1}$
@@
    Calculate $\mathbf{f}_{NC}$ based on new $\dot{m}^{\prime\prime}$, $C^{n+1}$ and $\mathbf{u}^{*}$
@@ 3.2 Calculation of Interfacial Mass Flux
Different approaches have been implemented in the literature to obtain $\dot{m}^{\prime\prime}$ from Eq. (8), focusing on the calculation of the temperature gradients in each phase, e.g., [8, 38, 10, 42, 13, 1, 17, 35]. Here, a normal stencil or probe is extended from the interface plane centroid into each phase along $\mathbf{n}_{\Gamma}$ [43, 44, 16]. Then, scalar quantities from the grid nodes, i.e., temperature, are linearly interpolated on two probing points on each side. To avoid using values from the opposite phase during interpolation, the first point has to be placed sufficiently far away from the interface. One can determine that the minimum distance to avoid this conflict is the cell diagonal (i.e., $\sqrt{2}\Delta x$ in two dimensions and $\sqrt{3}\Delta x$ in three dimensions). Consequently, the first point is placed at a distance of $\Delta x^{(1)}=1.75\Delta x$ from the interface. Additionally, a second point is placed at $\Delta x^{(2)}=\Delta x$ from the first point (i.e., $\Delta x^{(1)}+\Delta x^{(2)}=2.75\Delta x$). Although this is not problematic in the problems analyzed in this work, one may shorten the stencil and use a single point if it extends into a conflicting interface, e.g., two bubbles coming very close to each other.
A two-dimensional sketch is shown in Figure 2 for a non-conflicting scenario. One-sided second-order finite differences with non-uniform spacing can be used to calculate
@@
where, oftentimes, $T_{\Gamma}$ equals the saturation temperature at the given system pressure $T_{sat}$. An inherent issue of the calculation of $\nabla T_{G}\cdot\mathbf{n}_{\Gamma}$ and $\nabla T_{L}\cdot\mathbf{n}_{\Gamma}$ is its effect on the smoothness of the $\dot{m}^{\prime\prime}$ distribution. As such, some works perform weighted averages of the temperature gradients around the interface cell to obtain a more uniform distribution [10, 13, 17]. Despite a smoother $\dot{m}^{\prime\prime}$ distribution is observed with the proposed gradient calculation compared to other works for similar grid resolutions, variations in $\dot{m}^{\prime\prime}$ still affect the long-term evolution of canonical configurations such as static evaporating droplets or growing bubbles. Therefore, once Eq. (8) is evaluated, the value of $\dot{m}^{\prime\prime}$ is recalculated at each interface cell by averaging $\dot{m}^{\prime\prime}$ from interface cells contained in a stencil of size $3\times 3\times 3$. As shown in Section 5.1, even though the first probing point extends almost two grid sizes into each phase, the second-order one-sided scheme results in $\dot{m}^{\prime\prime}$ errors and convergence rates similar to other works.
@@ 3.3 Solution of the Navier-Stokes Equations
The Navier-Stokes equations are solved using a modification of the predictor-projection method [45] together with machine-accurate direct solvers for Poisson-type equations based on Fast Fourier Transforms (FFT) [46, 47], which have been incorporated in the solver using the vfftpack library [^2] A vectorized package of Fortran subprograms for the fast Fourier transform of multiple real sequences, R.A. Sweet, L.L. Lindgren and R.F. Boisvert, https://www.netlib.org/vfftpack/.
@@
Then, Eq. (21) is a Poisson equation for $\psi$ obtained by taking the divergence of Eq. (20), which is solved using the FFT-based or direct solver with the same boundary conditions as the pressure field. As shown later in Section 3.5, a pressure jump that does not correspond to the expected physical jump is introduced if this first predictor-projection step is not included (i.e., if $\mathbf{u}^{*}=\mathbf{u}^{n}$ remains true), resulting from the different Stefan flow from $t^{n}$ to $t^{n+1}$ and directly linked to the first term on the LHS of Eq. (12).
$\nabla^{2}\psi=\nabla\cdot\mathbf{u}^{n}-\nabla\cdot\mathbf{u}^{*}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n}-\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n+1}$ (21)
@@
with, e.g., $\nabla\psi_{i+\frac{1}{2}}\approx(\psi_{i+1}-\psi_{i})/\Delta x$ at the cell faces. Note some indices have been dropped for convenience. Here, $\delta_{\Gamma}=||\nabla C||$ is evaluated from staggered volume fractions to minimize the smearing of the velocity jump when compared to using central differences [48]. For example, Figure 1 shows how $C_{i+\frac{1}{2}}$ is obtained from adding the enclosed volumes on the right half of $C_{i}$ and on the left half of $C_{i+1}$. This evaluation of $\delta_{\Gamma}$ results in a more uniform and smoother distribution of the volumetric source term around curved interfaces compared to obtaining $\delta_{\Gamma}$ from the surface area density, e.g., [10, 16], especially at lower interface resolutions. This preserves the symmetries of the flow around, e.g., an evaporating droplet better while maintaining a reasonably sharp phase jump across a region of thickness $\sim 2\Delta x$.
Next, the second predictor step calculates a velocity $\mathbf{u}^{**}$ from the momentum equation, Eq. (15), without the pressure gradient, $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$ using the AB scheme
@@
As done for the energy equation, AB is used everywhere outside $\Omega_{\Gamma}$, while a classic first-order scheme is used in $\Omega_{\Gamma}$. The viscous term is discretized with second-order central differences, and the convective term is discretized with the third-order WENO scheme in cells not belonging to $\Omega_{\Gamma}$ and second-order central differences in $\Omega_{\Gamma}$. These modifications around the interface are necessary to aim for the cancellation of the numerical jump caused by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ with $\mathbf{f}_{NC}$. Note the choice of WENO in the momentum equation is to maintain consistency with the discretization of the energy equation. For practical purposes, no major differences have been observed between using WENO or central differences in the problems considered in this work.
The variable-coefficient pressure Poisson equation (PPE) resulting from the projection step is converted into a constant-coefficient PPE via the approximate substitution [49, 9]
@@
which splits the gradient $\tfrac{1}{\rho^{n+1}}\nabla p^{n+1}$ into a constant-coefficient implicit term and a variable-coefficient explicit term. For accuracy and stability purposes, $\rho_{0}=\text{min}(\rho_{g},\rho_{l})$ and $\hat{p}=[1+(\Delta t^{n+1}/\Delta t^{n})]p^{n}-(\Delta t^{n+1}/\Delta t^{n})p^{n-1}$ is a linear extrapolation in time of the pressure field [49, 9]. Then, a direct pressure solver can be used, resulting in a computationally more efficient solver than iterative methods [9, 50, 11, 12, 16].
However, pressure-splitting introduces considerable time-stepping restrictions for stability purposes when the pressure jump across the interface is large (e.g., large surface tension) or when the density ratio is large (i.e., $\nabla\hat{p}$ has a greater contribution in the PPE) [9]. In other words, smoothness in $\hat{p}$ may be lost and substantial pressure oscillations occur as the interface moves across grid cells [51, 52]. Such limitation could defy the implementation of the gradient splitting technique and reduce the benefits of a direct solver. This problem is mitigated by explicitly accounting for the pressure jump in the PPE [51], similar to a GFM. Thus, $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$ are removed from Eq. (15) during the predictor step in Eqs. (23) and (24) and, instead, are added in the projection step. Within the context of the gradient splitting technique, this is equivalent to the temporal extrapolation of the pressure jump instead of the pressure itself, i.e., see Eq. (26).
@@
$\mathbf{u}^{n+1}=\mathbf{u}^{**}-\Delta t^{n+1}\Bigg{[}\frac{1}{\rho_{0}}\bigg{(}\nabla p^{n+1}-\mathbf{f}_{\sigma}^{n+1}-\mathbf{f}_{\dot{m}^{\prime\prime}}^{n+1}\bigg{)}+\bigg{(}\frac{1}{\rho^{n+1}}-\frac{1}{\rho_{0}}\bigg{)}\bigg{(}\nabla\hat{p}-\hat{\mathbf{f}}_{\sigma}-\hat{\mathbf{f}}_{\dot{m}^{\prime\prime}}\bigg{)}-\frac{1}{\rho^{n+1}}\mathbf{f}_{NC}^{n+1}\Bigg{]}$ (26)
@@
$\nabla^{2}p^{n+1}=\nabla\cdot\bigg{[}\bigg{(}1-\frac{\rho_{0}}{\rho^{n+1}}\bigg{)}\bigg{(}\nabla\hat{p}-\hat{\mathbf{f}}_{\sigma}-\hat{\mathbf{f}}_{\dot{m}^{\prime\prime}}\bigg{)}+\mathbf{f}_{\sigma}^{n+1}+\mathbf{f}_{\dot{m}^{\prime\prime}}^{n+1}+\frac{\rho_{0}}{\rho^{n+1}}\mathbf{f}_{NC}^{n+1}\bigg{]}+\frac{\rho_{0}}{\Delta t^{n+1}}\big{[}\nabla\cdot\mathbf{u}^{**}-\nabla\cdot\mathbf{u}^{n+1}\big{]}$ (27)
@@
where $\nabla\cdot\mathbf{u}^{n+1}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}\big{|}^{n+1}$ ensures that the correct velocity jump given by Eq. (10) is imposed. It is important to highlight that the splitting operator is not applied to $\mathbf{f}_{NC}^{n+1}$ since the variable density $\rho^{n+1}$ effectively cancels with the density term inside $\mathbf{f}_{NC}^{n+1}$. Then, Eq. (27) is solved using the FFT-based or direct solver. Another advantage of using the direct solver relates to stiffness problems that might arise as $\nabla\cdot\mathbf{u}^{n+1}$ increases, which could affect the convergence rate of iterative solvers [8, 11].
@@
The forces $\mathbf{f}_{\sigma}$, $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$ must be calculated at cell faces to correspond with the evaluation of $\nabla p$. In $\mathbf{f}_{\sigma}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$, $\kappa$ and $\dot{m}^{\prime\prime}$ are averaged at cell faces [9] and $\mathbf{n}_{\Gamma}\delta_{\Gamma}=-\nabla C$ following the CSF approach. $\nabla C$ is evaluated from cell values, e.g., $\partial C/\partial x|_{i+1/2}\approx(C_{i+1}-C_{i})/\Delta x$. In contrast, $\mathbf{f}_{NC}$ is discretized differently for smoothness purposes in order to minimize geometrical errors that may arise from an ill-defined discrete force at the interface given the sensitivity of the term $\mathbf{u}^{*}\cdot\mathbf{n}_{\Gamma}$. Note that the corrective nature of $\mathbf{f}_{NC}$ to counteract the error introduced by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ in $\Omega_{\Gamma}$ implies that $\mathbf{u}^{*}$ substitutes $\mathbf{u}$ in Eq. (13). In this work, $\mathbf{u}^{*}$ is linearly averaged at cell faces while $\mathbf{n}_{\Gamma}$ is obtained at staggered interface cells (e.g., $10<C_{i+1/2}<$ 1) using the MYC method with staggered volume fractions. Then, a narrow band of staggered cells around the interface is populated with a weighted average of the normal unit vectors obtained in the previous step with the method described in A. Lastly, discretizing $\mathbf{n}_{\Gamma}\delta_{\Gamma}$ with $\delta_{\Gamma}$ obtained in staggered locations as, e.g., $\delta_{\Gamma,i+1/2}=\tfrac{1}{2}(\delta_{\Gamma,i}+\delta_{\Gamma,i+1})$ improves the smoothness of $\mathbf{f}_{NC}$ compared to the substitution $\mathbf{n}_{\Gamma}\delta_{\Gamma}=-\nabla C$. As shown in Section 3.5, both $\delta_{\Gamma,i+1/2}$ or $\nabla C$ evaluated with staggered central differences, e.g., $\partial C/\partial x|_{i+1/2}\approx(C_{i+3/2}-C_{i-1/2})/(2\Delta x)$, result in the correction of the error introduced by $\mathbf{u}^{*}\cdot\nabla\mathbf{u}^{*}$ at the discrete level for a well-posed one-dimensional case.
Lastly, it is worth discussing the computational cost of the flow solver, in particular the inclusion of additional Poisson-type equations. Besides the PPE, a Poisson equation is solved to calculate the divergence-free velocity for the advection of $C$ in Section 2.2 and to calculate $\mathbf{u}^{*}$ via Eqs. (20) and (21). Depending on the set tolerance and type of, e.g., multigrid iterative solver, direct solvers may provide a speed-up between 3 and 14 times in typical two-phase flow implementations [9, 53]. This translates into the PPE representing only about 58% of the wall-clock time per time step in parallel computing implementations of non-evaporative flows [9]. However, many configurations of interest involve disperse bubbly or droplet-laden flows and load balancing easily deteriorates when using typical domain decompositions, e.g., pencils. That is, only a few processes might take care of geometric reconstructions, evaluating interfacial heat and mass transfer, or calculating localized forces. As such, a behavior closer to what is reported in [11] is observed where solving each additional Poisson equation represents about 5% of the cost per time step in a configuration like the static evaporating droplet in Section 4.2.